# Consumption Partial Insurance

### v5

@@ 4.1 Data
As is evident from the previous section, panel data on income and consumption are needed to identify the moments of the permanent and transitory shocks and the degrees of insurance to them. The Panel Study of Income Dynamics (PSID) provides such data for a representative and continuously evolving sample of U.S. households since $1968$. For most of this period, the PSID collected expenditures only on food items. Since $1999$, however, the PSID was drastically redesigned in that it started collecting information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016). We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when we started working on this paper; the data are available only biennially. This is in contrast to BPP, who use PSID data over 19801980198019801992199219921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
@@
Our baseline sample consists of married couples in the PSID with a male spouse aged $30$ to $65$. We focus on a sample that is representative of the U.S. population, eliminating the supplementary low-income subsample. Households are also removed in the event of a break-up or remarriage; our analysis thus focuses on income risk rather than on divorce or other household dissolution factors. Naturally, we require non-missing data on income, expenditure, and basic demographics. [^19] We also remove observations in the bottom 0.25% of the distributions of $\Delta y_{it}\Delta y_{it+1}$ and $\Delta c_{it}\Delta c_{it+1}$, i.e., large increases followed by large drops, likely reflecting measurement error, and observations with extreme jumps in income or consumption (jump: $\geq$10-fold increase or $\geq$90reduction), which likely also reflect measurement error. Our results are not sensitive to this sample selection choice. Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
@@
Appendix table C.1 presents summary statistics over the sample period 19991999199919992019201920192019. About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education. Our measure of income is household disposable income, namely the sum of taxable income (e.g., earnings and asset income), transfers, and social security income of the couple and of other family members, minus taxes paid. Average disposable income is \$142,910 in 2018 prices. [^20] We express all monetary figures in 2018 prices because earnings are 1-year retrospective and our latest data are from 2019. We deflate using the Consumer Price Index from the Bureau of Labor Statistics. Our measure of consumption is real expenditure on nondurable goods and services, comprising food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses (gasoline, parking, insurance), education, and child care. Average real consumption is \$26,766, with food at home accounting for about $30\%$ of this. Housing, including rent and home insurance, is the main expenditure we exclude; we do so for comparability with BPP who also exclude it. Including housing does not meaningfully change the subsequent results.
@@ 4.3 Empirical moments of income and consumption
Income moments. Panel A presents the moments of income growth. The variance is $0.153$, while the first-order autocovariance is $-0.049$ and highly significant. The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data. [^26] The second-order autocovariance is an order of magnitude smaller and only marginally significant. This feature and the biennial nature of the data, whereby the second-order autocovariance reflects correlations 4 calendar years apart, motivates our white noise specification for the transitory shock.
@@
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.049 (0.002) -0.049 (0.002)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.143 (0.098) -0.143 (0.098)
@@
${\operatorname*{\mathrm{Cov}}}(\Delta c_{it},\Delta c_{it+1})$0.044 (0.001) -0.140 (0.009)
${\operatorname*{\mathrm{Skew}}}(\Delta c_{it})$0.039 (0.034) -0.177 (0.083)
${\operatorname*{\mathrm{Cov}}}((\Delta c_{it})^{2},\Delta c_{it+1})$0.003 (0.001) 0.190 (0.036)
@@
The third standardized moment of income growth is $-0.143$. Despite the small size of the cross-section and the inherent difficulties in estimating higher moments, the point estimate misses statistical significance at the 10% level only by a small margin. ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$, which reflects (minus) the third central moment of the transitory shock, is significant at the 5% level. The point estimates suggest that income growth exhibits left skewness, i.e., a long left tail. In simple terms, the left tail of the distribution of income growth (i.e., of income shocks) is longer than the right tail, so a negative shock is on average more severe, more unsettling, than a positive one. This is in line with Guvenen et al. (2021), who document strong left skewness in the distribution of male earnings growth using administrative data from the U.S. Social Security Administration.
@@
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate a comparison, appendix C.2 reports moments of male earnings growth in our sample. The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively. Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al. (2021) also confirmed this. Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). [^27] Using earnings as the relevant income variable in the model does not materially change our subsequent results, in particular the asymmetry in the transmission of bad versus good (earnings) shocks. These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
@@
Consumption moments. Panel B of table 1 presents the moments of consumption growth; column 1 specifically reports moments from our baseline sample in the PSID. The variance of consumption growth is similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error. The first-order autocovariance is $-0.044$ and highly significant. In the context of our model, this autocovariance reflects the variance of consumption error $\sigma_{u^{c}_{t}}^{2}$ (see appendix B).
@@
The moments of imputed consumption growth are in column 2 of table 1. The variance at $0.323$ is two-and-a-half times larger than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The magnitude of the higher-order moments is also multiple times larger in the imputed versus the true series. Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
@@ 5.1 Income process
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at $-0.947$ and $-1.31$ for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. These results vary with the level of income, but we postpone that discussion for later. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are farther away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
@@
Seen together, large left skewness and excess kurtosis imply that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad. Ignoring these stark features of income shocks may lead to erroneous inference of the degree of consumption partial insurance, to which we now turn.
@@ 5.3 Quadratic consumption function
We now turn to the estimation of the quadratic consumption function. The results appear in column 5 of table 3; the full set of targeted moments is shown in appendix B.4.3.373737Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 2., [^38] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case, i.e., in column 4, table 3. This is because in the quadratic case certain 3rd- and 4th-order consumption moments require up to 8th moments of income shocks, which we do not model. This raises an issue of comparability between the linear and quadratic specifications, as the target moments differ (appendix table B.2). However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway. Our estimates of the linear function in that case do not differ from those in column 4, table 3, so estimation is robust to the choice of moments. These results are available upon request.
@@ 5.3.1 Main results
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of the consumption polynomial) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks but a positive $\psi^{(2)}=0.015$ for transitory ones. $\phi^{(2)}$ is statistically insignificant at the margin for now but turns highly significant subsequently. By contrast, $\psi^{(2)}$ is almost always insignificant, barring the few exceptions we highlight below. The coefficient $\omega^{(22)}$ on the product of the two shocks is large but always imprecise. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^39] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
@@
First, bad (negative) permanent shocks are more unsettling than good (positive) ones because they are associated with a larger transmission parameter into consumption. Consider, for example, a large permanent shock to household income, $|\zeta_{it}|=0.5$, about three standard deviations away from the mean. If the shock is negative, $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$ (s.e.$=0.036$, estimated by the delta method); by contrast, if the shock is positive, then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$ (s.e.$=0.036$). We can generalize this to bad and good shocks of any magnitude and conclude that
@@
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Compare, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$). We can generalize this to bad shocks of any magnitude and conclude that
@@
Why would a positive or small shock be better insured? When a good shock hits, households save part of it and consumption does not respond too much. When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements. Yet there is clearly a limit to how much they can dissave or borrow. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption. A large negative shock thus affects consumption by relatively more than a small or positive shock does. [^40] The results also imply that the pass-through of a good shock decreases with its magnitude. The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks. However, large left skewness and excess kurtosis imply that extreme good shocks are rare. We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
@@
$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.\$
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption.

### v6

@@ 4.1 Data
As is evident from the previous section, panel data on income and consumption are needed to identify the moments of the permanent and transitory shocks and the degrees of insurance to them. The Panel Study of Income Dynamics (PSID) provides such data for a representative and continuously evolving sample of U.S. households since $1968$. For most of this period, the PSID collected expenditures only on food items. Since $1999$, however, the PSID was drastically redesigned to collect information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016). We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when we started working on this paper; the data are available only biennially. This is in contrast to BPP, who use PSID data over 1980198019921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
@@
Our baseline sample consists of married couples in the PSID with a male spouse aged $30$ to $65$. We focus on a sample that is representative of the U.S. population, eliminating the supplementary low-income subsample. Households are also removed in the event of a break-up or remarriage; our analysis thus focuses on income risk rather than on divorce or other household dissolution factors. Naturally, we require non-missing data on income, expenditure, and basic demographics. [^19] We also remove observations in the bottom 0.25% of the distributions of $\Delta y_{it}\Delta y_{it+1}$ and $\Delta c_{it}\Delta c_{it+1}$, i.e., large increases followed by large drops, likely reflecting measurement error, and observations with extreme jumps in income or consumption (jump: $\geq$10-fold increase or $\geq$90% reduction), which likely also reflect measurement error. Our results are not sensitive to this sample selection choice. Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
@@
Appendix table C.1 presents summary statistics over the sample period 1999199920192019. About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education. Our measure of income is household disposable income, namely the sum of taxable income (e.g., earnings and asset income), transfers, and social security income of the couple and of other family members, minus taxes paid. Average disposable income is \$142,910 in 2018 prices. [^20] We express all monetary figures in 2018 prices because earnings are 1-year retrospective and our latest data are from 2019. We deflate using the Consumer Price Index from the Bureau of Labor Statistics. Our measure of consumption is real expenditure on nondurable goods and services, comprising food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses (gasoline, parking, insurance), education, and child care. Average real consumption is \$26,766, with food at home accounting for about $30\%$ of this. Housing, including rent and home insurance, is the main expenditure we exclude; we do so for comparability with BPP who also exclude it. Including housing does not meaningfully change the subsequent results.
@@ 4.3 Empirical moments of income and consumption
Income moments. Panel A presents the moments of income growth. The variance is $0.153$, while the first-order autocovariance is 0.049-0.049 and highly significant. The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data. [^26] The second-order autocovariance is an order of magnitude smaller and only marginally significant. This feature and the biennial nature of the data, whereby the second-order autocovariance reflects correlations 4 calendar years apart, motivates our white noise specification for the transitory shock.
@@
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$ -0.049 (0.002) -0.049 (0.002)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$ -0.143 (0.098) -0.143 (0.098)
@@
${\operatorname*{\mathrm{Cov}}}(\Delta c_{it},\Delta c_{it+1})$ -0.044 (0.001) -0.140 (0.009)
${\operatorname*{\mathrm{Skew}}}(\Delta c_{it})$ -0.039 (0.034) -0.177 (0.083)
${\operatorname*{\mathrm{Cov}}}((\Delta c_{it})^{2},\Delta c_{it+1})$ -0.003 (0.001) 0.190 (0.036)
@@
The third standardized moment of income growth is 0.143-0.143. Despite the small size of the cross-section and the inherent difficulties in estimating higher moments, the point estimate misses statistical significance at the 10% level only by a small margin. ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$, which reflects (minus) the third central moment of the transitory shock, is significant at the 1% level. The point estimates suggest that income growth exhibits left skewness, i.e., a long left tail. In simple terms, the left tail of the distribution of income growth (i.e., of income shocks) is longer than the right tail, so a negative shock is on average more severe, more unsettling, than a positive one. This is in line with Guvenen et al. (2021), who document strong left skewness in the distribution of male earnings growth using administrative data from the U.S. Social Security Administration.
@@
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate a comparison, appendix C.2 reports moments of male earnings growth in our sample. The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively. Skewness is 0.569-0.569 (and highly statistically significant) versus 1.039-1.039 in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID. Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
Evidently, left skewness and excess kurtosis are not only features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). [^27] Using earnings as the relevant income variable in the model does not materially change our subsequent results, in particular the asymmetry in the transmission of bad versus good (earnings) shocks. These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
@@
Consumption moments. Panel B of table 1 presents the moments of consumption growth; column 1 specifically reports moments from our baseline sample in the PSID. The variance of consumption growth is similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error. The first-order autocovariance is 0.044-0.044 and highly significant. In the context of our model, this autocovariance reflects the variance of consumption error $\sigma_{u^{c}_{t}}^{2}$ (see appendix B).
@@
The moments of imputed consumption growth are in column 2 of table 1. The variance at $0.323$ is two-and-a-half times larger than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The magnitude of the higher-order moments is also multiple times larger in the imputed versus the true series. Skewness in imputed consumption growth is 0.177-0.177 (versus 0.039-0.039 in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
@@ 5.1 Income process
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at 0.947-0.947 and 1.31-1.31 for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. These results vary with the level of income, but we postpone that discussion for later. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are farther away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
@@
Seen together, left skewness and excess kurtosis imply that most households face either negligible shocks (close to zero) or shocks that are quite bad. Ignoring these stark features may lead to erroneous inference of the degree of partial insurance, to which we now turn.
@@ 5.3 Quadratic consumption function
We now turn to the estimation of the quadratic consumption function. The results appear in column 5 of table 3; the full set of targeted moments is shown in appendix B.4.3.373737Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 2.
@@ 5.3.1 Main results
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of the consumption polynomial) at $\phi^{(1)}=0.134$, close to the values 0.110.110.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks but a positive $\psi^{(2)}=0.015$ for transitory ones. $\phi^{(2)}$ is statistically insignificant at the margin for now but turns highly significant subsequently. By contrast, $\psi^{(2)}$ is almost always insignificant. The coefficient $\omega^{(22)}$ on the product of the two shocks is large but always imprecise. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^38] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income., [^39] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case in column 4, table 3. This is because in the quadratic case certain 3rd- and 4th-order consumption moments require up to 8th moments of income shocks, which we do not model. This raises an issue of comparability between the linear and quadratic specifications, as the target moments differ (appendix table B.2). However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway. Our estimates of the linear function in that case do not differ from those in column 4, table 3, so estimation is robust to the choice of moments. These results are available upon request.
@@
First, bad (negative) permanent shocks are more unsettling than good (positive) ones because they are associated with a larger transmission parameter into consumption. Consider, for example, a large permanent shock to household income, $|\zeta_{it}|=0.5$, about three standard deviations away from the mean. If the shock is negative, $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$ (s.e.=0.036=0.036, estimated by the delta method); by contrast, if the shock is positive, then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$ (s.e.=0.036=0.036). We can generalize this to bad and good shocks of any magnitude and conclude that
@@
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Compare, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.=0.044=0.044) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.=0.026=0.026). We can generalize this to bad shocks of any magnitude and conclude that
@@
Why would a positive or small shock be better insured? When a good shock hits, households save part of it and consumption does not respond too much. When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements. Yet there is clearly a limit to how much they can dissave or borrow. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption. A large negative shock thus affects consumption by relatively more than a small or positive shock does. [^40] The results also imply that the pass-through of a good shock decreases with its magnitude. The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks. However, large left skewness and excess kurtosis imply that large good shocks are rare. We should thus not extrapolate the implications of the parameter estimates for the case of large good shocks, as such shocks are effectively not present in our sample.
@@
$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.\\$
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption.